105. Calc Air drag is a significant problem in som
Answer #1

a(v)=g-\beta v^{2}\Rightarrow \frac{\mathrm{d} v}{\mathrm{d} t}=g-\beta v^{2}=\beta (\frac{g}{\beta }-v^{2})

\Rightarrow \frac{dv}{(\frac{g}{\beta }-v^{2})}=\beta dt

Let (g/B)=k2.

Now ,\frac{1}{(k^{2}-v^{2})}=\frac{1}{2k}(\frac{1}{k+v}+\frac{1}{k-v})

So,the differential equation becomes

\frac{1}{2k}(\frac{1}{k+v}+\frac{1}{k-v})dv=\beta dt

Integrate this equation with limits :

  • v = 0 to v = v
  • t = 0 to t = t

to get   \frac{1}{2k}[\ln (k+v)-\ln (k-v)]=\beta t

\Rightarrow ln(\frac{k+v}{k-v})=2k\beta t

\Rightarrow (\frac{k+v}{k-v})=e^{2k\beta t}

\Rightarrow v=k [\frac{e^{2k\beta t}-1}{e^{2k\beta t}+1}]\Rightarrow v=k [\frac{e^{2k\beta t}-1}{e^{2k\beta t}+1}]= k \tanh (k\beta t)

Substitute k=sqrt(g/B) to get

\Rightarrow v= \sqrt{\frac{g}{\beta }} \tanh (\sqrt{(g\beta )} t)

The terminal velocity is that at which acceleration 'a' is zero and and object moves uniformly with this velocity.

\Rightarrow a_{terminal}=g-\beta v_{terminal}^{2}=0

\Rightarrow v_{terminal}=\sqrt{\frac{g}{\beta }}